OpenStudy (anonymous):

Solve for x: −3|x + 7| = −12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (anonymous):

First step is to get the absolutes alone. So divide that -3 over to the other side.

OpenStudy (anonymous):

\[\left| x \right|=a,x= \pm a\]

OpenStudy (anonymous):

\[\left| x-a \right|=b,x-a= \pm b\]

OpenStudy (anonymous):

I keep trying these but I never understand them.

OpenStudy (anonymous):

divide by -3 and start with \[|x+7|=4\] then solve \[x+7=4\] and \[x+7=-7\] separately

OpenStudy (anonymous):

both are easy to solve, subtract 7 from both sides

OpenStudy (anonymous):

dang i meant \[x+7=4\\ x+7=-4\]both are solved by subtracting 7 from both sides

OpenStudy (anonymous):

WOW I ACTUALLY GOT IT THANK YOU

OpenStudy (anonymous):

Wait I think I got it wrong, I got x = −3, x = 11